# Borsuk-Ulam and Replicable Learning of Large-Margin Halfspaces

### v3

@@ Abstract
Recent remarkable advances in learning theory have established that, for total concept classes, list replicability, global stability, differentially private (DP) learnability, and shared-randomness replicability all coincide with the finiteness of Littlestone dimension. Does this equivalence extend to partial concept classes?
We answer this question by proving that the list replicability number of $d$-dimensional $\gamma$-margin half-spaces satisfies
@@
which grows with dimension. Consequently, for partial classes, list replicability and global stability do not necessarily follow from bounded Littlestone dimension, pure DP-learnability, or shared-randomness replicability.
Applying our main theorem, we resolve several open problems:
@@
  • There exists a partial concept class with Littlestone dimension $1$ such that all its disambiguations have infinite Littlestone dimension. This answers a question of Cheung, H. Hatami, P. Hatami, and Hosseini (ICALP ’23).
Our lower bound follows from a topological argument based on the local Borsuk-Ulam theorem of Chase, Chornomaz, Moran, and Yehudayoff (STOC ’24). For the upper bound, we construct a list-replicable learning rule using the generalization properties of SVMs.
@@ 1 Introduction
The large-margin half-space classification problem is a central topic in learning theory, extensively studied for its theoretical significance and practical impact. Theoretically, it is a fundamental and elegant geometric problem, and it serves as a simple model for understanding more complex learning problems. Practically, its importance stems from the success of Support Vector Machines (SVMs), which leverage the large-margin assumption to produce accurate classifications in high-dimensional spaces. SVMs play a key role in many real-world applications, including pattern and text recognition, image analysis, geographic information systems, bioinformatics, healthcare, and fraud detection.
To describe the large-margin classification problem, we first consider standard half-space classification without any margin assumption. After appropriate normalization and homogenization, we may assume that the domain is the unit sphere $\mathbb{S}^{d-1}\subseteq\mathbb{R}^{d}$, and the objective is to learn an unknown homogeneous half-space which classifies each point on the sphere as either ±$\pm 1$ depending on whether it belongs to the half-space.
The standard mathematical framework for analyzing the complexity of a learning task is probably approximately correct (PAC) learning. In this setting, the learner receives a training set of labeled points drawn independently from an unknown distribution $\mathcal{D}$, and the goal is to produce a hypothesis that accurately predicts the labels of new points sampled from $\mathcal{D}$ with high probability. The fundamental theorem of PAC learning states that the size of the training set required for PAC learning depends on a combinatorial parameter known as the VC dimension (Definition 2.5).
We also consider the Littlestone dimension (Definition 2.7), a refinement of the VC dimension that determines the optimal number of mistakes in online learning. In online learning, the learner receives data points sequentially and must predict the label of each point in turn before observing its correct label. The goal is to minimize the number of mistakes throughout the learning process.
Since the VC dimension of homogeneous half-spaces in $\mathbb{R}^{d}$ is $d$, PAC-learning this class requires a large training set in the high-dimensional settings. Moreover, the Littlestone dimension of this class is infinite except in the trivial case of $d=1$, meaning the class is not online learnable, even in $\mathbb{R}^{2}$.
Under the large-margin assumption, the learning task is restricted to points that lie at least a margin $\gamma>0$ away from the defining hyperplane. This constraint allows algorithms such as Perceptron [MP43, Ros58] and Support Vector Machines [VC64, CV95] to achieve efficient PAC and online learning, even in high-dimensional spaces. We study the large-margin halfspaces problem through the formal lens of partial classes, which offers a general framework for analyzing such constrained learning tasks.
@@
A partial concept class over an arbitrary domain $\mathcal{X}$ is a set {±$\mathcal{H}\subseteq\left\{\pm 1,\star\right\}^{\mathcal{X}}$, where each $h\in\mathcal{H}$ is called a partial concept. The value $h(x)=\star$ indicates that $h$ is undefined at $x$, and therefore, both ±$\pm 1$ are acceptable predictions for the label of $x$.
@@
In the large-margin setting, the domain is $\mathbb{S}^{d-1}$ and every homogeneous half-space defines a partial concept that assigns $h(x)=\star$ if $x$ lies within distance $\gamma$ of the defining hyperplane of $h$. Otherwise, it classifies $x$ as ±$\pm 1$ depending on whether it belongs to the half-space. More formally, each hypothesis $h_{w}:\mathbb{S}^{d-1}\to\{\pm 1,\star\}$ is specified by a unit vector $w\in\mathbb{S}^{d-1}$ and given by
@@
$h_{w}(x)\coloneqq\begin{cases}\operatorname{sgn}(\langle w,x\rangle)&\text{if }|\langle w,x\rangle|\geq\gamma\\ \star&\text{otherwise}\end{cases}.$ (1)
We denote the partial concept class of all such $h_{w}$ by $\mathcal{H}^{d}_{\gamma}$.
We also define $\mathcal{H}^{\infty}_{\gamma}$ as the class of all partial concepts $h_{w}:\bigcup_{d\in\mathbb{N}}\mathbb{S}^{d-1}\to\left\{\pm 1,\star\right\}$, where each hypothesis is specified by a unit vector $w\in\bigcup_{d\in\mathbb{N}}\mathbb{S}^{d-1}$ of arbitrary finite dimension. For any $x$ and $w$ of different dimensions, we set $h_{w}(x)\coloneqq\star$, and otherwise, define $h_{w}(x)$ as in Eq. 1.
@@
In [AHHM22], Alon, Hanneke, Holzman, and Moran proved that, as in the case of total concept classes, the VC and Littlestone dimensions characterize PAC and online learnability of partial concept classes. In the case of $\mathcal{H}_{\gamma}^{d}$, the classic mistake-bound analysis of the Perceptron algorithm [MP43, Ros58] (see also [SSBD14, Theorem 9.1]) shows the following upper bound on the Littlestone and VC dimensions:
@@
Recent advances in learning theory, sparked by the influential works on the differential privacy of PAC learning [ALMM19, BLM20, ABL+22], have established that for total concept classes, finite Littlestone dimension not only characterizes online learnability but is also equivalent to list replicability, global stability, differential privacy (DP) learnability, and shared-randomness replicability. We will discuss these notions in depth shortly. This naturally leads to the question of whether these equivalences extend to partial concept classes. The primary goal of this paper is to provide a negative answer to this question by proving that $\mathcal{H}_{\gamma}^{\infty}$ is not list replicable, even though it has a bounded Littlestone dimension by Eq. 2.
Before delving deeper into these results, we formally define some key notions. Consider a partial concept class {±$\mathcal{H}\subseteq\{\pm 1,\star\}^{\mathcal{X}}$. A distribution $\mathcal{D}$ on $\mathcal{X}\times\left\{\pm 1\right\}$ is realizable by a partial concept class {±$\mathcal{H}\subseteq\left\{\pm 1,\star\right\}^{\mathcal{X}}$ if, for every $n$, a random sample $\bm{S}=((\bm{x}_{i},\bm{y}_{i}))_{i=1}^{n}\sim\mathcal{D}^{n}$ is almost surely realizable by some $h\in\mathcal{H}$, meaning that $h(\bm{x}_{i})=\bm{y}_{i}$ for all $i=1,\ldots,n$. Note that $\mathcal{D}$ is a distribution over $\mathcal{X}\times\left\{\pm 1\right\}$, so none of the labels $\bm{y}_{i}$ take the value $\star$.
Here, and throughout the paper, we use boldface letters to denote random variables and use the notation $(\bm{x},\bm{y})\sim\mathcal{D}$ to express that $(\bm{x},\bm{y})$ is a random variable distributed according to $\mathcal{D}$.
The population loss of a partial concept $h\in\left\{\pm 1,\star\right\}^{\mathcal{X}}$ with respect to a distribution $\mathcal{D}$ on $\mathcal{X}\times\left\{\pm 1\right\}$ is
$\mathcal{L}_{\mathcal{D}}(h)\coloneqq\Pr_{(\bm{x},\bm{y})\sim\mathcal{D}}[h(\bm{x})\neq\bm{y}].$
Replicability.
Replicability refers to the ability of an algorithm or study to produce consistent results when repeated under similar conditions and with similar data. It is a key principle of the scientific method that ensures research findings are reliable and not due to chance.
In recent years, a rapidly growing body of research has emerged that introduced various rigorous formulations of replicability for learning algorithms, particularly in the context of the PAC learning framework [BLM20, MM22, CMY23, BGH+23, KVYZ23, EKK+23, EKM+23, MSS23, EHKS23, KKL+24, KKMV23]. We introduce only the notions that are central to our main results and refer the reader to prior works, such as [MSS23, KKMV23, MM22], for a discussion of other related notions. Throughout, we use replicability as an umbrella term referring to the various formulations of this general concept.
A learning rule $\bm{\mathcal{A}}$ for a concept class {±$\mathcal{H}\subseteq\left\{\pm 1,\star\right\}^{\mathcal{X}}$ is $(\epsilon,\rho)$-globally stable if for every realizable distribution $\mathcal{D}$, there is a hypothesis $h\in\left\{\pm 1\right\}^{\mathcal{X}}$ with population loss $\mathcal{L}_{\mathcal{D}}(h)\leq\epsilon$ satisfying
$\Pr_{\bm{S}\sim\mathcal{D}^{n}}[\bm{\mathcal{A}}(\bm{S})=h]\geq\rho,\ \ \text{where }n=n(\epsilon).$
In other words, when run with samples from $\mathcal{D}$, there is a non-negligible chance $\rho>0$ that the learner will output the same hypothesis $h$. We define $\rho^{\mathrm{gs}}_{\epsilon}(\mathcal{H})$ to be the supremum of $\rho$ such that there is a $(\epsilon,\rho)$-globally stable learner for $\mathcal{H}$. The global stability parameter of $\mathcal{H}$ is then defined as
$\rho^{\mathrm{gs}}(\mathcal{H})\coloneqq\inf_{\epsilon>0}\rho^{\mathrm{gs}}_{\epsilon}(\mathcal{H}).$
The definition of global stability might initially seem weak, as a globally stable learner is not necessarily a PAC learner. In particular, since $\rho$ can be a small constant, there may be a probability as great as $1-\rho$ that the learning rule outputs a hypothesis with large population loss. However, [CMY23] showed that for total classes {±$\mathcal{H}\subseteq\left\{\pm 1\right\}^{\mathcal{X}}$, global stability is equivalent to a stronger notion called list replicability.
Definition 1.1 (List replicability).
A learning rule $\bm{\mathcal{A}}$ with sample complexity $n(\epsilon,\delta)$ is said to be an $(\epsilon,L)$-list replicable learner for {±$\mathcal{H}\subseteq\{\pm 1,\star\}^{\mathcal{X}}$ if the following holds. For every realizable distribution $\mathcal{D}$ on $\mathcal{X}\times\left\{\pm 1\right\}$, there is a list of total hypotheses $h_{1},\ldots,h_{L}\in\left\{\pm 1\right\}^{\mathcal{X}}$ such that
  • The population loss of every $h_{i}$ with respect to $\mathcal{D}$ is at most $\epsilon$:
    $\mathcal{L}_{\mathcal{D}}(h_{i})\leq\epsilon\text{ for every }i=1,\ldots,L.$
  • The probability that $\bm{\mathcal{A}}$ outputs a hypothesis outside the list is at most $\delta$:
    $\Pr_{\bm{S}\sim\mathcal{D}^{n}}[\bm{\mathcal{A}}(\bm{S})\not\in\left\{h_{1},\ldots,h_{L}\right\}]\leq\delta\ \text{ where }n=n(\epsilon,\delta).$
Given an accuracy parameter $\epsilon\in(0,1)$, the $\epsilon$-list replicability number of $\mathcal{H}$, denoted by $\operatorname{LR}_{\epsilon}(\mathcal{H})$, is the smallest integer $L$, such that there exists an $(\epsilon,L)$-list replicable learner for $\mathcal{H}$. We define $\operatorname{LR}_{\epsilon}(\mathcal{H})=\infty$ if no such integer exists.
The list replicability number of $\mathcal{H}$ is $\operatorname{LR}(\mathcal{H})\coloneqq\sup_{\epsilon>0}\operatorname{LR}_{\epsilon}(\mathcal{H}).$ We say $\mathcal{H}$ is list replicable if $\operatorname{LR}(\mathcal{H})<\infty$.
In [CMY23], Chase, Moran and Yehudayoff proved that for every total class {±$\mathcal{H}\subseteq\left\{\pm 1\right\}^{\mathcal{X}}$, list replicability is equivalent to global stability. It is easy to check that their proof applies to partial concept classes, resulting in the following relationship between the list replicability number and global stability parameter.
Theorem 1.2 ([CMY23]).
For every partial concept class {±$\mathcal{H}\subseteq\left\{\pm 1,\star\right\}^{\mathcal{X}}$, we have
$\rho^{\mathrm{gs}}(\mathcal{H})=\frac{1}{\operatorname{LR}(\mathcal{H})}.$
For completeness, we provide a proof in Appendix A.
As shown in [BLM20], the global stability parameter of any nontrivial concept class is at most $1/2$. This fundamental limitation implies that no nontrivial class can achieve true replicability in the strongest sense, where the learning algorithm outputs the same hypothesis with high probability. To remedy this, Impagliazzo, Lei, Pitassi, and Sorrell [ILPS22] introduced a weaker notion of replicability where different executions of the algorithm can use the same random seed. Formally, let $\mathcal{A}(S,r)$ be a randomized learning rule, where $r$ denotes the random seed.
Definition 1.3 (Shared-randomness replicability).
A concept class {±,}$\mathcal{H}\subseteq\left\{\pm,\star\right\}^{\mathcal{X}}$ is shared-randomness replicable if there exists a learning rule $\mathcal{A}$ and a sample complexity function $n(\epsilon,\delta)$ such that, for every $\epsilon,\delta>0$ and every realizable distribution $\mathcal{D}$, the following conditions hold:
  • Small population loss: $\Pr_{\bm{S}\sim\mathcal{D}^{n},\bm{r}}[\mathcal{L}_{\mathcal{D}}\left(\mathcal{A}(\bm{S},\bm{r})\right)>\epsilon]\leq\delta.$
  • Replicability with shared randomness: $\Pr_{\bm{S,S^{\prime}}\sim\mathcal{D}^{n},\bm{r}}[\mathcal{A}(\bm{S},\bm{r})=\mathcal{A}(\bm{S^{\prime}},\bm{r})]\geq 1-\delta.$
The following theorem, which brings together several celebrated recent results in learning theory, establishes that for total concept classes, all these notions of replicability, along with approximate differential privacy, are characterized by the finiteness of the Littlestone dimension.
Theorem 1.4 ([ALMM19, BLM20, ABL+22, CMY23, ILPS22]).
Let {±$\mathcal{H}\subseteq\left\{\pm 1\right\}^{\mathcal{X}}$ be a total concept class. The following statements are equivalent.
  • $\operatorname{Ldim}(\mathcal{H})<\infty$.
  • $\mathcal{H}$ is globally stable and equivalently list replicable.
  • $\mathcal{H}$ is shared-randomness replicable.
  • $\mathcal{H}$ is approximately differentially private (DP)-learnable (See Definition 2.9).
1.1 Main contributions
Replicability of large-margin half-spaces.
The differentially private and replicable learning of large-margin half-spaces have been studied extensively [BDMN05, LNUZ20, BMNS19, KMST20, BCS20, BMS22a, BMS22b, ILPS22, KKL+24].
In [LNUZ20], Nguyen, Ullman, and Zakynthinou gave a DP-learner for $\mathcal{H}^{d}_{\gamma}$ with a sample complexity that does not depend on the dimension $d$. In fact, they proved a stronger result that $\mathcal{H}^{d}_{\gamma}$ is pure DP-learnable (Definition 2.10). Combining the dimension-free DP-learnability of $\mathcal{H}^{d}_{\gamma}$ with the “DP-learnable to Shared-randomness Replicability” reduction of [BGH+23] shows that $\mathcal{H}^{d}_{\gamma}$ is shared-randomness replicable with dimension-free sample complexity. More recently, Kalavasis, Karbasi, Larsen, Velegkas, and Zhou [KKL+24] employed a more sophisticated rounding scheme to directly obtain a shared-randomness replicability of this problem, which further improves the sample complexity. In the following theorem, we summarize the consequences of these lines of work that are relevant to this paper.
Theorem 1.5 ([MP43, Ros58, LNUZ20, KKL+24]).
For every $\gamma>0$ and $d\in\mathbb{N}\cup\left\{\infty\right\}$, the class $\mathcal{H}^{d}_{\gamma}$ satisfies the following.
  • $\operatorname{Ldim}(\mathcal{H}^{d}_{\gamma})<\gamma^{-2}$.
  • The class $\mathcal{H}^{d}_{\gamma}$ is (pure) DP-learnable with dimension-independent sample complexity.
  • The class $\mathcal{H}^{d}_{\gamma}$ is shared-randomness replicable with dimension-independent sample complexity.
Given the positive results in Theorem 1.5 and the equivalences in Theorem 1.4, one might expect that, as in the case of total concepts, the list replicability number of $\mathcal{H}^{d}_{\gamma}$ is independent of $d$. Our main theorem establishes that this is not the case—the list replicability number of $\mathcal{H}^{d}_{\gamma}$ grows as $d$ increases.

### v4

@@ Abstract
We prove that the list replicability number of $d$-dimensional $\gamma$-margin half-spaces satisfies
@@
which grows with dimension. This resolves several open problems:
@@
  • There is a separation of $O(1)$ vs $\omega(1)$ between randomized and pseudo-deterministic communication complexity.
@@
  • There exists a partial concept class with Littlestone dimension $1$ such that all its disambiguations have infinite Littlestone dimension. This resolves a problem of Cheung, H. Hatami, P. Hatami, and Hosseini (ICALP ’23).
Our lower bound follows from a topological argument based on a local Borsuk-Ulam theorem. For the upper bound, we construct a list-replicable learning rule using the generalization properties of SVMs.
@@ 1 Introduction
Large-margin half-space classification is a fundamental problem in learning theory. In this setting, data is normalized to lie on the unit sphere $\mathbb{S}^{d-1}\subset\mathbb{R}^{d}$, and we are guaranteed a promise that each point lies at least a fixed margin $\gamma\in(0,1)$ from some unknown homogeneous hyperplane. The learner is then tasked with classifying points based on the side of the defining hyperplane to which they belong. This problem has been extensively studied for both its theoretical and practical significance: it provides a clean geometric model for analyzing more complex learning tasks, and underlies the success of Support Vector Machines (SVMs), which leverage the large-margin assumption to produce accurate classifications in high-dimensional spaces, with applications across domains such as text and image recognition, bioinformatics, and fraud detection.
We study this problem through the lens of replicability, the requirement that an algorithm produce consistent outcomes when repeated under similar conditions and with similar data. In recent years, replicability has become a vibrant research area, and various rigorous formulations of replicability for learning algorithms have been introduced and studied [BLM20, MM22, CMY23, BGH+23, KVYZ23, EKK+23, EKM+23, MSS23, EHKS23, KKL+24, KKMV23]. Among them, one of the most intriguing is the notion of global stability, which was first discovered in connection with differentially private learning and online learning [BLM20, ABL+22, CMY23]. Subsequent work, however, has shown that its significance extends well beyond these applications. In its equivalent formulation as list replicability [CMY23], the notion is intrinsically linked to the geometry and topology of the space of realizable distributions of a concept class [CMY23, CCMY24, BGHH25, CMW25, BHH+25]. For example, for finite classes, it characterizes the topological dimension of this space under its natural simplicial structure [BHH+25]. Through these connections, fundamental results in classical topological dimension theory, such as the Lebesgue covering theorem, translate directly into statements about learnability and replicability.
Our main result states that the list replicability number of the large-margin classification problem in $\mathbb{R}^{d}$ lies between $\tfrac{d}{2}+1$ and $d$. In particular, it diverges as the ambient dimension $d$ grows. This stands in contrast to many common complexity measures for the same task, such as the VC dimension, Littlestone dimension, and randomized communication complexity, which are bounded by a function of the margin $\gamma\in(0,1)$ independent of the ambient dimension $d$.
This divergence has several consequences and resolves a number of open problems from previous works, as discussed in detail in Section 2.1. Beyond its implications within learning theory, it also connects naturally to questions in communication complexity. In particular, our most notable consequence shows that any disambiguation of the large-margin classification problem into a total concept class must have a large Littlestone dimension and a large randomized communication complexity. Thus, while the original partial problem is “easy” under these classical measures, every possible completion of it to a total problem is inherently “hard”. Establishing such lower bounds for disambiguations is typically very challenging, as the initial partial problem is “easy” and one has no control over how it is extended to a total one.
A notable consequence of our disambiguation theorem is an $O(1)$ versus $\Omega(\log\log n)$ gap between randomized and pseudo-deterministic communication complexities. Separating these two measures is a well-known open problem in the study of pseudodeterminism [GIPS21], and our result provides the first $O(1)$ versus $\omega(1)$ separation.
1.1 Preliminaries
We study the large-margin half-space problem through the formal lens of partial classes, which offers a general framework for analyzing such constrained learning tasks.
@@ 1.1 Preliminaries
A partial concept class over an arbitrary domain $\mathcal{X}$ is a set $\mathcal{C}\subseteq\left\{\pm 1,\star\right\}^{\mathcal{X}}$, where each $c\in\mathcal{C}$ is called a partial concept. The value $c(x)=\star$ indicates that $c$ is undefined at $x$, and therefore, both $\pm 1$ are acceptable predictions for the label of $x$.
PAC learning.
The standard mathematical framework for analyzing the complexity of a learning task is probably approximately correct (PAC) learning. In PAC learning, the learner is given parameters $\delta,\epsilon>0$ and receives training data consisting of $n=n(\mathcal{C},\delta,\epsilon)$ independent labeled examples drawn from an unknown but fixed distribution $\mu$ over $\mathcal{X}\times\{\pm 1\}$. We work in the realizable setting: for every $n$, a random sample $\bm{S}=((\bm{x}_{i},\bm{y}_{i}))_{i=1}^{n}\sim\mu^{n}$ is almost surely realizable by some $c\in\mathcal{C}$, meaning that $c(\bm{x}_{i})=\bm{y}_{i}$ for all $i=1,\ldots,n$. [^1] Here, and throughout the paper, we use boldface letters to denote random variables and use the notation $(\bm{x},\bm{y})\sim\mu$ to express that $(\bm{x},\bm{y})$ is a random variable distributed according to $\mu$. Note that $\mu$ is a distribution over $\mathcal{X}\times\left\{\pm 1\right\}$, so none of the labels $\bm{y}_{i}$ take the value $\star$. The learner’s task is to use the training data to output, with probability at least $1-\delta$, a hypothesis $h:\mathcal{X}\to\{\pm 1\}$ whose population loss
$\operatorname{loss}_{\mu}(h)\coloneqq\Pr_{(\bm{x},\bm{y})\sim\mu}[h(\bm{x})\neq\bm{y}]$
is at most $\epsilon$.
The following simple lemma from [AHHM21] establishes the connection between realizability and having zero population loss.
Lemma 1.1 ([AHHM21]).
Let $\mathcal{C}\subseteq\left\{\pm 1,\star\right\}^{\mathcal{X}}$ be a partial concept class, and let $\mu$ be a distribution on $\mathcal{X}\times\left\{\pm 1\right\}$. If $\operatorname{loss}_{\mu}(\mathcal{C})\coloneqq\inf_{c\in\mathcal{C}}\operatorname{loss}_{\mu}(c)$ is zero, then $\mu$ is realizable by $\mathcal{C}$. Conversely, if $\mu$ is realizable and has finite or countable support, then $\operatorname{loss}_{\mu}(\mathcal{C})=0$.
The fundamental theorem of PAC learning states that the size of the training set required for PAC learning a total concept class depends on a combinatorial parameter known as the VC dimension, which we now define in the more general partial setting. A (partial) concept class $\mathcal{C}\subseteq\{\pm 1,\star\}^{\mathcal{X}}$ shatters a set $S\subseteq\mathcal{X}$ if $\{c|_{S}:c\in\mathcal{C}\}=\{\pm 1\}^{S}$, where $c|_{S}$ denotes the restriction of $c$ to $S$. The VC dimension of $\mathcal{C}$ is defined as
$\operatorname{VCdim}(\mathcal{C})\coloneqq\sup\{|S|:\ S\subseteq\mathcal{X}\text{ is shattered by }\mathcal{C}\}.$
In [AHHM21], Alon, Hanneke, Holzman, and Moran proved that the fundamental theorem of PAC learning holds for partial concept classes as well.
List replicability.
Definition 1.2 (List replicability).
A learning rule $\bm{\mathcal{A}}$ is an $(\epsilon,L)$-list replicable learner for $\mathcal{C}\subseteq\left\{\pm 1,\star\right\}^{\mathcal{X}}$ if for every $\delta>0$, there is a sample complexity $n=n(\delta)$ such that the following holds. For every realizable distribution $\mu$ on $\mathcal{X}\times\left\{\pm 1\right\}$, there exists $h_{1},\ldots,h_{L}\in\left\{\pm 1\right\}^{\mathcal{X}}$ such that
$\operatorname{loss}_{\mu}(h_{i})\leq\epsilon~\forall i~\text{ and }~\Pr_{\bm{S}\sim\mu^{n}}[\bm{\mathcal{A}}(\bm{S})\not\in\left\{h_{1},\ldots,h_{L}\right\}]\leq\delta.$
The $\epsilon$-list replicability number of $\mathcal{C}$ is
$\operatorname{LR}_{\epsilon}(\mathcal{C})\coloneqq\min\{L:\exists(\epsilon,L)\text{-list replicable learner for }\mathcal{H}\},$
with $\operatorname{LR}_{\epsilon}(\mathcal{C})=\infty$ if none exists. The list replicability number of $\mathcal{C}$ is
$\operatorname{LR}(\mathcal{C})\coloneqq\sup_{\epsilon>0}\operatorname{LR}_{\epsilon}(\mathcal{C}).$
We say $\mathcal{C}$ is list replicable if $\operatorname{LR}(\mathcal{C})<\infty$.
Remark 1.3.
Some readers might be familiar with an equivalent form of list replicability known as global stability. Its definition is not needed for the main results of this paper, so we include it in Section A.3, along with the definition of the global stability parameter $\rho(\mathcal{C})$ of a concept class $\mathcal{C}$. This parameter is analogous to the list replicability number, and in fact [CMY23] proved that the two quantities are related by the equation $\rho(\mathcal{C})=1/\operatorname{LR}(\mathcal{C})$ for total concept classes. It is easy to check that the proof extends to the partial setting (see Appendix B). Hence, qualitative results about list replicability also hold for global stability, and quantitative results hold after the appropriate reciprocal modification.
Online learning and Littlestone dimension.
In online learning, a learner receives data points sequentially from an adversary and must predict each label before seeing the correct answer. The goal is to minimize the total number of mistakes. The optimal mistake bound is captured by the Littlestone dimension, a refinement of the VC dimension defined via mistake trees.
A mistake tree of depth $d$ over domain $\mathcal{X}$ is a complete binary tree whose internal nodes are labeled by points $x\in\mathcal{X}$ and edges by bits $b\in\{\pm 1\}$ (1-1 for left, +1+1 for right). Following a path from the root to a leaf thus yields a sequence $(x_{1},b_{1}),\dots,(x_{d},b_{d})$, where each $x_{i}$ is the node label at level $i$ and $b_{i}$ records whether the path goes left or right.
A concept class $\mathcal{C}\subseteq\{\pm 1,\star\}^{\mathcal{X}}$ shatters such a tree if for every root-to-leaf path there exists $c\in\mathcal{C}$ with $c(x_{i})=b_{i}$ for all $i$. The Littlestone dimension $\operatorname{Ldim}(\mathcal{C})$ is the largest $d$ for which some depth-$d$ mistake tree is shattered, or $\infty$ if no such $d$ exists.
It always holds that $\operatorname{VCdim}(\mathcal{C})\leq\operatorname{Ldim}(\mathcal{C})$, since any set $S=\left\{x_{1},\ldots,x_{d}\right\}$ shattered by $\mathcal{C}$ gives rise to a mistake tree of depth $d$, where all nodes at level $i$ are labeled with $x_{i}$. This tree is shattered by $\mathcal{C}$.
Littlestone proved that a total concept class $\mathcal{C}$ is online learnable if and only if $\operatorname{Ldim}(\mathcal{C})<\infty$. This result was later extended to partial concept classes by [AHHM21]
Large-margin half-spaces.
@@
In the large-margin setting, the domain is $\mathbb{S}^{d-1}$ and every homogeneous half-space defines a partial concept that assigns $c(x)=\star$ if $x$ lies within distance $\gamma$ of the defining hyperplane of $h$. Otherwise, it classifies $x$ as $\pm 1$ depending on whether it belongs to the half-space. More formally, each concept $c_{w}:\mathbb{S}^{d-1}\to\{\pm 1,\star\}$ is specified by a unit vector $w\in\mathbb{S}^{d-1}$ and given by
@@
$c_{w}(x)\coloneqq\begin{cases}\operatorname{sgn}({\left\langle w,x\right\rangle})&\text{if }|{\left\langle w,x\right\rangle}|\geq\gamma\\ \star&\text{otherwise}\end{cases}.$ (1)
We denote the partial concept class of all such $c_{w}$ by $\mathcal{H}^{d}_{\gamma}$.
For the standard half-space classification problem $\mathcal{H}^{d}$ without any margin assumption (that is, each $x$ is labeled by $\operatorname{sgn}({\left\langle w,x\right\rangle})$ whenever ${\left\langle w,x\right\rangle}\neq 0$ and by $\star$ otherwise), we have $\operatorname{VCdim}(\mathcal{H}^{d})=d$. Moreover, $\operatorname{Ldim}(\mathcal{H}^{d})=\infty$, except in the trivial case of $d=1$. In particular, this class is not online learnable, even in $\mathbb{R}^{2}$.
@@
However, under the large-margin assumption $\gamma>0$, the classic mistake-bound analysis of the Perceptron algorithm [MP43, Ros58] (see also [SSBD14, Theorem 9.1]) shows the following upper bound on the Littlestone and VC dimensions:
@@
Gap Hamming problem.
$\operatorname{GHD}_{\gamma}^{n}(x,y)\coloneqq\begin{cases}\operatorname{sgn}({\left\langle x,y\right\rangle})&\text{if $|{\left\langle x,y\right\rangle}|>\gamma n$}\\ \star&\text{otherwise}\end{cases}.$
As a communication problem, Alice receives $x$, Bob receives $y$, and under the promise $|{\left\langle x,y\right\rangle}|\geq\gamma n$, they must compute $\operatorname{GHD}_{\gamma}^{n}(x,y)$ with minimal communication. For fixed $\gamma$, the public-coin randomized communication complexity of $\operatorname{GHD}_{\gamma}^{n}$ is $O_{\gamma}(1)$: using shared randomness, the players sample a subset $S\subseteq[n]$ of size $O_{\gamma}(1)$ and estimate $\left\langle x,y\right\rangle$} by $\tfrac{n}{|S|}\sum_{i\in S}x_{i}y_{i}$, which requires only $2|S|$ communicated bits.
2 Main theorem
Our main theorem determines the list replicability number of $\mathcal{H}_{\gamma}^{d}$ up to a factor of two, showing in particular that it grows unboundedly with the dimension $d$.